QUANTUM TRANSLATION-ROTATION DYNAMICS OF HYDROGEN MOLECULES CONFINED IN THE CAGES OF CLATHRATE HYDRATES

\underline{ZLATKO BAČIĆ}, FRANCESCO SEBASTIANELLI, MINZHONG XU, YAEL S. ELMATAD and JULES W. MOSKOWITZ, Department of Chemistry, New York University, New York, NY 10003.

The coupled translation-rotation (T-R) eigenstates of a hydrogen molecule inside the small dodecahedral (H_{2}O)_{20} cage of the structure II clathrate hydrate have been determined accurately by means of quantum 5D calculations, for para- and ortho-H_{2}^{a}, as well as ortho- and para-D_{2}^{b}. In addition, the ground-state properties of two and three para-H_{2} and ortho-D_{2} molecules confined in the small cage have been calculated rigorously using the diffusion Monte Carlo method^{b}. These calculations have provided a comprehensive picture of the quantum T-R dynamics of the encapsulated molecules. The translational modes exhibit negative anharmonicity; j is a good rotational quantum number, with the threefold degeneracy of the j = 1 level lifted completely. When two hydrogen molecules are confined, they are effectively excluded from the central region of the cage, and reside within a shell less than 2 bohrs wide. If time permits, the quantum dynamics results for multiple H_{2}/D_{2} molecules inside the large (H_{2}O)_{28} cage will be presented.

aM. Xu, Y. S. Elmatad, F. Sebastianelli, J. W. Moskowitz, and Z. Bačić, J. Phys. Chem. B \underline{110}, 24806 (2006)

bF. Sebastianelli, M. Xu, Y. S. Elmatad, J. W. Moskowitz, and Z. Bačić, J. Phys. Chem. C \underline{111}, 2497 (2007)